OpenStudy (anonymous):

Write an equation in slope-intercept form of the line that passes through the point (-4, 1) and has a slope of -(3/2)

jimthompson5910 (jim_thompson5910):

y = mx+b 1 = (-3/2)(-4)+b 1 = 12/2+b 1 = 6+b 1 - 6 = b -5 = b b = -5 Answer is \[\Large y = -\frac{3}{2}x-5\]